### v2

# Optimizing Retinal Prosthetic Stimuli with Conditional Invertible Neural Networks

Abstract
Implantable retinal prostheses offer a promising solution to restore partial vision by circumventing damaged photoreceptor cells in the retina and directly stimulating the remaining functional retinal cells. However, the information transmission between the camera and retinal cells is often limited by the low resolution of the electrode array and the lack of specificity for different ganglion cell types, resulting in suboptimal stimulations. In this work, we propose to utilize normalizing flow-based conditional invertible neural networks to optimize retinal implant stimulation in an unsupervised manner. The inherent invertibility of these networks allows us to use them as a surrogate for the computational model of the visual system, while also encoding input camera signals into optimized electrical stimuli on the electrode array. Compared to other methods, such as trivial downsampling, linear models, and feed-forward convolutional neural networks, the flow-based invertible neural network and its conditional extension yield better visual reconstruction qualities w.r.t. various metrics using a physiologically validated simulation tool.
1 Introduction
Figure 1: (a) Anatomy of the retina with an electrical epiretinal implant. Modified based on Retinal Implant by Mbuerki under CC BY-SA 3.0 [12]. (b) Electrode activation with the axon map model (top, by Beyeler et al. [8] under CC BY 4.0) and a retinal implant with a $9\times 9$ electrode grid on the axon map (bottom, generated with pulse2percept [7] ).
Retinal implants [60, 65] aim to restore partial vision to individuals with certain types of visual impairments, particularly those caused by diseases like retinitis pigmentosa (RP) or age-related macular degeneration (AMD) [33, 61, 48, 25, 42]. The visual signals captured by an external camera are transmitted to the electrical signals on the implanted electrode array, which bypass damaged photoreceptor cells in the retina and directly stimulate the remaining functional retinal cells (Figure 1(a)). Various approaches exist to implant the prosthesis to different locations, such as epiretinal (facing the ganglion cells), subretinal (behind the photoreceptors), and suprachoroidal (on top of the choroid). Since we utilize the Axon Map Model (Figure 1(b)) as the simulation model from the stimulus to the percept, which is based on the epiretinal implant Argus I/II devices [35, 10], we limit our findings for epiretinal implants [56], such as Argus I/II [35, 10], EPIRET3 [38, 45], and VLARS [59, 34]. However, the limited resolution and lack of specificity for different ganglion cell types (such as ON and OFF cells) necessitate an advanced algorithm for stimulation optimization [47, 23, 28, 36]. The retinal prosthetic stimulation optimization can be considered as an inverse problem, where the optimized stimulus is achieved with the inverse transformation of the biologically fixed visual pathway (Figure 2). With the help of powerful machine learning and deep learning techniques [9, 5], the optimization tasks can be solved, e.g., using an evolutionary heuristic algorithm [46], a reconstruction decoder [55] or reinforcement learning [31]. Based on a physiologically validated simulation tool, pulse2percept [7], a line of deep learning-based approaches are proposed to learn an inverse model [43, 63], to learn a hybrid autoencoder [19], or to learn an optimization encoder with a surrogate visual pathway simulation model in an end-to-end manner [62]. Besides, preference-based human-in-the-loop Bayesian optimization algorithms are proposed [16, 18], which facilitate determining the patient-individual parameters of an in silico retinal prosthetic computational model.
As one of the notable early papers of Invertible Neural Networks (INNs), Dinh et al. [14] propose to use INNs for density estimation tasks. Throughout a series of non-volume transformation processes, the normalizing flow [44] (total probability mass) is preserved, which ensures that the resulting distribution remains a valid probability distribution. Conceptually, INNs are naturally suitable to construct an autoencoder [30], as the forward and backward direction of the INN should form a perfect encoder-decoder pair (but not always [6]). Nguyen et al. [40] train an INN autoencoder with an artificial bottleneck, created by setting several elements to zero before the inverse pass. Sorrenson and Draxler et al. [50] train autoencoders with likelihood on novel flow architectures that do not rely on coupling and are not dimension preserving. Furthermore, Ardizzone et al. first propose an INN variant to solve inverse problems with the maximum mean discrepancy loss (INN-MMD) [3] and an extension to INNs with explicit controls, namely Conditional Invertible Neural Networks (cINNs) [4, 2]. Analogous to other generative models, cINNs can also be applied to unsupervised data synthesis, such as [15]. Zhou et al. [64] apply flow-based invertible generative models to neural encoding and decoding, transforming natural images and corresponding neural spike trains recorded from salamander retinal ganglion cells (RGC) [41] bidirectionally. One key methodological distinction in our approach is that we directly train the stimulus-percept pairs using a physiologically validated simulation tool [7]. In contrast, the neural signals (spike trains) in [64] are trained to reconstruct visual signals without a strong constraint in the decoder in the salamander RGC dataset [41].
Figure 2: A schematic overview of the stimulation optimization for retinal prostheses. The visual signals are converted directly into electrical signals on the electrodes through an optimization process, constrained by a simulation model of the visual pathway.
In this work, we follow an INN [3] and a cINN approach [4, 2] to learn a computational model for the visual pathway (Figure 2) and automatically obtain the inverse mapping from percept to stimulus thanks to the intrinsic feature of invertibility. Furthermore, INNs are generative and can be implemented to become bijective to solve inverse problems [3]. In addition, by imposing conditioning to the INN, the generated inverse input (optimized stimulus, see Figure 3) shall follow the desired constraint. We compare the reconstruction quality of the predicted percepts using cINNs to other approaches, including trivial downsampling, a linear model, feed-forward and invertible neural networks, and address the promising optimization for retinal prosthetic stimuli.
2 Basics of Invertible Neural Network
Given a random variable $\mathbf{z}=h\left(\mathbf{x}\right)$ with a known and tractable density, e.g., a spherical multivariate normal distribution $\mathbf{z}\sim\pi\left(\mathbf{z}\right)=\mathcal{N}\left(\mathbf{z};\mathbf{0},\mathbf{I}\right)$, a data point $\mathbf{x}$ from an i.i.d. dataset $\mathcal{D}$ can be generated by $\mathbf{x}=h^{-1}(\mathbf{z})$ for an invertible function $h$. Following the change of variables theorem, the probability density function of such a model given a data point $\mathbf{x}$ can be written as:
$p(\mathbf{x})=\pi(h(\mathbf{x}))\left|\det\frac{\partial h(\mathbf{x})}{\partial\mathbf{x}^{T}}\right|=\pi(\mathbf{z})\left|\det\mathbf{J}_{h}(\mathbf{x})\right|,$ (1)
where $\mathbf{J}_{h}(\mathbf{x})=\partial h(\mathbf{x})/\partial\mathbf{x}^{T}$ is the Jacobian matrix of $h$ at $\mathbf{x}$ and $\left|\det\mathbf{J}_{h}(\mathbf{x})\right|$ is the absolute value of the Jacobian determinant. Different names describe this approach from different perspectives of the likelihood (Equation 1): Invertible neural network (INN) points out the intrinsic architecture design of $h$ shall be invertible; Normalizing flow (NF) emphasizes that a data distribution of $\mathbf{x}$ is normalized into a simpler form of $\mathbf{z}$ by a series of transformations (flow); Non-volume preserving (NVP, [14]) specifies that the scaling factor $\left|\det\mathbf{J}_{h}(\mathbf{x})\right|$ of volume elements under the flow transformation is not constrained to be $1$.
With $\mathbf{x}$ and $\mathbf{z}$ sharing the identical dimensionality $\mathbf{x},\mathbf{z}\in\mathbb{R}^{n}$ (cf. Section 3.2), the normalizing flow $h$ can be trained by maximizing likelihood of $\mathbf{x}$, which is equivalent to minimizing the Negative Log-Likelihood (NLL) loss:
$\mathcal{L}_{\mathrm{NLL}}$ $=-\log(p(\mathbf{x}))$ (2)
$=-\log\left(\pi(\mathbf{z})\right)-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|$ (3)
$=-\log\left(\mathcal{N}\left(\mathbf{z};\mathbf{0},\mathbf{I}\right)\right)-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|$ (4)
$=-\log\prod_{i=1}^{n}\left(\frac{1}{\sqrt{2\pi}}e^{-\frac{z_{i}^{2}}{2}}\right)-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|$ (5)
$=\frac{n}{2}\log(2\pi)+\frac{1}{2}\|\mathbf{z}\|_{2}^{2}-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|$ (6)
$\simeq\frac{1}{2}\|h\left(\mathbf{x}\right)\|_{2}^{2}-\log\left|\det\mathbf{J}_{h}\left(\mathbf{x}\right)\right|.$ (7)
The first term in Equation 7 can be easily calculated with the output of the forward normalizing flow. The second term is computationally expensive mainly due to the matrix determinant of the Jacobian. A line of INN architecture designs, e.g., additive coupling [13], affine coupling [14] and Glow coupling [29], has been proposed to utilize an LU decomposition style architecture (Equation 13) to achieve the following properties. (1) Invertibility: The design of the coupling layer ensures that the transformation is invertible, which is crucial for computing the exact likelihood of the data under the transformed distribution. (2) Computational efficiency: Both forward and inverse transformations are computationally efficient, which allows the model to scale to high-dimensional data. (3) Easy Jacobian: The Jacobian matrix of the transformation is triangular, making the computation of its determinant straightforward. The coupling layer is structurally similar to the lifting scheme architecture [51] used in discrete wavelet transform. In an affine coupling layer [14], the input is split into two parts with $\mathbf{x}:=[\mathbf{x}_{A},\mathbf{x}_{B}]$, only one part of which is applied with an affine transformation and the other remains unchanged:
$\begin{cases}\mathbf{z}^{\prime}_{A}=\mathbf{x}_{A}\\ \mathbf{z}^{\prime}_{B}=\mathbf{x}_{B}\odot\exp(s(\mathbf{x}_{A}))+t(\mathbf{x}_{A})\,,\end{cases}$ (8)
where $\odot$ denotes the element-wise multiplication, while $s(\cdot)$ and $t(\cdot)$ are scale and translation functions, respectively. $\mathbf{z}^{\prime}:=[\mathbf{z}^{\prime}_{A},\mathbf{z}^{\prime}_{B}]$ represents the intermediate output of the forward one-sided transformation. The invertibility of the first half of the affine coupling layer is straightforward and efficient due to its structure with:
$\begin{cases}\mathbf{x}_{A}=\mathbf{z}^{\prime}_{A}\\ \mathbf{x}_{B}=\left(\mathbf{z}^{\prime}_{B}-t(\mathbf{x}_{A})\right)\odot\exp(-s(\mathbf{x}_{A}))\,.\end{cases}$ (9)
Note that $s(\cdot)$ and $t(\cdot)$ are not required to be invertible themselves and can be any differentiable function (to facilitate gradient-based optimization) including neural networks. The Jacobian matrix of this forward one-sided coupling transformation from $\mathbf{x}$ to $\mathbf{z}^{\prime}$ is triangular, making its log-determinant easy to compute:
$\log\left|\det(\mathbf{J})\right|$ $=\log\left|{\begin{array}[]{cccc}\mathbf{I}&\mathbf{0}\\ \frac{\partial\mathbf{z}^{\prime}_{B}}{\partial{\mathbf{x}_{A}}}&\text{diag}\left({\exp\left({s\left(\mathbf{x}_{A}\right)}\right)}\right)\\ \end{array}}\right|$ (12)
$=\sum s({\mathbf{x}_{A}}).$ (13)
Additionally, in the subsequent layer, the second segment remains unchanged, whereas the first segment will be transformed with the intermediate output from the other side:
$\begin{cases}\mathbf{z}_{A}=\mathbf{z}^{\prime}_{A}\odot\exp(s(\mathbf{z}^{\prime}_{B}))+t(\mathbf{z}^{\prime}_{B})\\ \mathbf{z}_{B}=\mathbf{z}^{\prime}_{B}\,.\end{cases}$ (14)
More details about the INN architectures used in this paper can be found in Section 4.3 and in Figure 4.
The original generative paradigm (e.g., RealNVP [14]) is not suitable for the application of retinal stimulus optimization, as the optimized stimulus shall be guided by the target visual signal as the reference. Therefore, we utilize two INN variants, namely INN-MMD [3] with the combination of an unsupervised MMD loss (Section 3.2) and a supervised MSE loss, and cINN [4, 2] with an explicit conditioning mechanism (Section 3.3).
(a) INN-MMD
(b) cINN
Figure 3: Training and evaluation pipelines for the stimulation optimization with (a) invertible neural networks with maximum mean discrepancy loss (INN-MMD) [3, 4] and (b) conditional invertible neural networks (cINN) [2].
3 Methods
3.1 Stimulation Optimization
We begin by formulating the in silico optimization problem for retinal prosthetic stimuli. Let ${f}$ represent the function of the visual pathway (Figure 2), which uses the electrical stimulus $\mathbf{x}_{s}$ from the implanted electrode array as input and produces the percept $\mathbf{y}$ as output. We assume that $f_{\phi}$ can simulate ${f}$, where $\phi$ denotes the parameters of the computational model, such as the Axon Map Model [8]. The desired encoding function $g_{\theta}:\tilde{\mathbf{x}}\mapsto\mathbf{x}_{s}$ from the visual signal (target) $\tilde{\mathbf{x}}$ to the stimulus ${\mathbf{x}_{s}}$ can be found with the following optimization problem:
[…]
where $\mathcal{L}_{\mathrm{eval}}$ serves as the objective function. We employ a neural network to construct $g_{\theta}$ with learnable parameters $\theta$. In practice, we often train a differentiable neural network as a surrogate for $f_{\phi}$ (e.g., [62, 19]), enabling convenient end-to-end optimization for the encoder $g_{\theta}$.
Figure 4: We follow [4, 2] to design the INN and cINN architectures with Glow coupling layers [29]. The conditional input $\mathbf{c}$ in both forward and inverse Glow coupling blocks is the only difference between the INN and the cINN approach.
3.2 Invertible Neural Network for Inverse Problems with Maximum Mean Discrepancy
If a process $q:\mathbf{x}_{D}\in\mathbb{R}^{D}\mapsto\mathbf{y}_{d}\in\mathbb{R}^{d}$ is not bijective with $D>d$, we can introduce a latent random variable $\mathbf{z}_{K}\sim\pi\left(\mathbf{z}_{K}\right)=\mathcal{N}\left(\mathbf{z}_{K};\mathbf{0},\mathbf{I}_{K}\right)$ with $K=D-d$, as proposed by Ardizzone et al. [3]. We then construct an invertible process $h:\mathbf{x}_{D}\mapsto[\mathbf{y}_{d},\mathbf{z}_{K}]$ with $h(\mathbf{x}_{D})=[h_{\mathbf{y}_{d}}(\mathbf{x}_{D}),h_{\mathbf{z}_{K}}(\mathbf{x}_{D})]$ and $h_{\mathbf{y}_{d}}(\mathbf{x}_{D})\approx q(\mathbf{x}_{D})$. Similar to [3], we apply a supervised Mean Squared Error (MSE) loss on $\mathbf{y}_{d}$ with $\mathcal{L}_{\mathrm{MSE}}=\mathbb{E}[(\mathbf{y}_{d}-h_{\mathbf{y}_{d}}(\mathbf{x}_{D}))^{2}]$ and an unsupervised Maximum Mean Discrepancy (MMD) loss [22, 21] in terms of $\mathbf{x}_{D}$ and $\mathbf{z}_{K}$, respectively. The MMD loss enforces two distributions $p_{M}$ and $p_{M^{\prime}}$ to be identical and is formulated as:
$\begin{split}\mathcal{L}_{\mathrm{MMD}}(p_{M},p_{M^{\prime}})=&\\ \left(\mathbb{E}_{i,j}[k(\mu_{i},\mu_{j})\right.&\left.-2k(\mu_{i},\mu_{j}^{\prime})+k(\mu_{i}^{\prime},\mu_{j}^{\prime})]\right)^{\frac{1}{2}},\end{split}$ (16)
where $\mathbb{E}$ denotes the expected value and $\{\mu\},\{\mu^{\prime}\}$ are two sets of samples according to the distributions $\mu\sim p_{M}$ and $\mu^{\prime}\sim p_{M^{\prime}}$, respectively.
As suggested in [3, 53], we utilize the radial basis function-based (RBF) rational quadratic kernel [39] (also known as inverse multiquadric kernel):
$k(\mu,\mu^{\prime})=\left(1+\left\|(\mu-\mu^{\prime})/\sigma\right\|_{2}^{2}\right)^{-1}\,.$ (17)
The MMD loss on $\mathbf{x}_{D}$ forces it to match the prior with $\mathcal{L}_{\mathrm{MMD}}(p(\mathbf{x}_{D}),p(h^{-1}([\mathbf{y}_{d},\mathbf{z}_{K}])))$ and is optional. The MMD loss on $\mathbf{z}_{K}$ forces it to match the normal distribution with $\mathcal{L}_{\mathrm{MMD}}(p(h(\mathbf{x}_{D})),p(\mathbf{y}_{d})p(\mathbf{z}_{K}))$. Moreover, following the approach in [3], we pad $\mathbf{x}_{D}$ with zeros to match the required dimensionality when necessary. This allows us to train both directions as the forward pass, as illustrated in Figure 5. Under the constraint of the invertible network architecture, both $h$ and $h^{-1}$ can be trained in both directions with a guaranteed invertible relation. Numerous invertible neural network architecture designs have been proposed to ensure that the inverse and the Jacobian determinant are easily computed, such as additive coupling layer [13], affine coupling layer [14] and Glow coupling layer [29] (see Section 2). The implementation details of the architecture in this work are visualized in Figure 4.
Table 1: Summary of notations for retinal implants (Section 3.1) in the context of invertible neural networks.
Notations INN-MMD [3] cINN [4, 2]
$f_{\phi}$ Forward network ($h$)
$g_{\theta}$ Reverse network ($h^{-1}$)
$\mathbf{x}_{s}$ Forward input ($\mathbf{x}$ or $\mathbf{x}_{D}$)
$\mathbf{\tilde{x}}$ Reverse input Reverse conditioning
$\mathbf{y}$, $\mathbf{y}_{d}$ Forward output (MSE) Forward conditioning
$\mathbf{z}$, $\mathbf{z}_{K}$ Forward output (MMD) Forward output (NLL)
3.3 Conditional Invertible Neural Network
Within the local coupling blocks [14], conditioning features $\mathbf{c}$ can be integrated into both directions of the network by concatenating them in the subfunctions with $\mathbf{z}=h(\mathbf{x};\mathbf{c})$ and $\mathbf{x}=h^{-1}(\mathbf{z};\mathbf{c})$, forming Conditional Invertible Neural Networks (cINNs) [4, 2]. This is similar to other conditional generative models, e.g., conditional variational auto-encoder (cVAE) [49] and conditional generative adversarial network (cGAN) [27]. Accordingly, the NLL loss for cINN maximum likelihood training is formulated as
$\mathcal{L}_{\mathrm{NLL}}\simeq\frac{1}{2}\|h(\mathbf{x};\mathbf{c})\|_{2}^{2}-\log\left|\det\mathbf{J}_{h}(\mathbf{x})\right|\,.$
As shown in Figure 3(b), we train the cINN by minimizing the NLL loss in an unsupervised manner and use the predicted percept during the training and the MNIST target during the evaluation as the conditioning. The same invertible neural network architecture is used as in the INN experiments. Unlike INN-MMD, which is constrained by a hybrid loss (combining supervised MSE and unsupervised MMD), cINN explicitly uses the conditioning as a reference to guide the stimulus with the target. A summary of notations for retinal implants from Section 3.1 is listed in Table 1, along with their corresponding contexts in INN-MMD and cINN.
(a) Forward process: percept-to-stimulus.
(b) Forward process: stimulus-to-percept.
Figure 5: MSE loss values of the INN-MMD training [3] (a) for percept-to-stimulus being the forward process and (b) for stimulus-to-percept being the forward process. We observe that the MSE loss values of inverse processes are always larger than the supervised forward processes, which implies that the invertibility using a supervised INN is not guaranteed. We report the results of the setup (a) in Table 2, as the optimized stimulus is desired.
4 Experiments and Results
4.1 Axon Map Model
In our in silico experiments, we create a virtual square retinal implant with a $9\times 9$ electrode grid (Figure 1(b)) and utilize the physiologically validated Axon Map Model [8], implemented in pulse2percept [7]. The phosphenes elicited by the axons, somas, or dendrites of the retinal ganglion cells have different shapes, as demonstrated in 1999 by Greenberg et al. [20]. The recent Axon Map Model [8] aligns with these findings by modelling the drawings from the subjects with retinal implants, which is one of the available models in pulse2percept [7] along with such as Scoreboard Model [8] and Biphasic Axon Map Model [17]. The predicted percept intensity at location $(x,y)$ is calculated by:
$\begin{split}I(x,y)=\exp&\left(-\frac{(x-x_{\text{stim}})^{2}+(y-y_{\text{stim}})^{2}}{2\rho^{2}}\right.\\ &\left.-\frac{(x-x_{\text{soma}})^{2}+(y-y_{\text{soma}})^{2}}{2\lambda^{2}}\right),\end{split}$ (18)
with exponential decays in terms of the distances to the stimulus at $(x_{\text{stim}},y_{\text{stim}})$ and to the elongated soma at $(x_{\text{soma}},y_{\text{soma}})$, controlled by the constants $\rho$ and $\lambda$, respectively (see Figure 1(b)). In this paper, we set the constants to the realistic and challenging values $\rho=400\,\mu m$ and $\lambda=1550\,\mu m$ for all experiments (cf. [8]).
4.2 Datasets
In order to achieve generalizability, models are trained using 10 million random stimulus-percept pairs and then evaluated with the MNIST test set [32]. We conduct all experiments using a batch size of 1024. The stimulus always has a resolution $9\times 9$ pixels (i.e., the number of the electrodes) and we experiment with both $9\times 9$ and $28\times 28$ pixels for percepts and the MNIST targets. Each stimulus-percept pair of the training set is generated by uniformly sampling in $[-3,+3]$ for each electrode. We select this intensity range to amplify the MSE values for particularly inaccurate predictions, when the difference between the predicted and ground-truth values exceeds 1. Negative voltage values in this range are beneficial in inhibiting phosphene elongation, which effectively designates the corresponding electrodes as return electrodes, as described in [54, 57]. We visualize stimuli and percepts with negative values in Figure 7, which do not result in any stimulated phosphenes though (similar to the neuron’s refractory period). Additionally, for each stimulus, only a random percentage of electrodes are activated while the rest are set to zero. Examples of stimulus-percept pairs are showcased in Figure 7(a).
(a) Correlation between log Jacobian determinant and MSE for MNIST and random samples.
(b) Improved MSE with the optimized $\mathbf{z}$ w.r.t. log Jacobian determinant for the 50 worst MNIST samples.
Figure 6: Investigation of likelihood (using log Jacobian determinant as in Equation 7) and MSE loss with cINNs. The y-axis is plotted on a logarithmic scale for better visualization.
4.3 Implementation Details and Preliminary Remarks
We compare five different approaches to optimize retinal prosthetic stimuli. The visualization of the invertible neural network architectures are depicted in Figure 4.
Downsampling. Trivial downsampling with a single learnable gain factor is applied to an input MNIST image as the visual signal.
Linear model. A single linear layer is used to learn the inverse transformation from percepts to stimuli. During the inference, the linear model outputs the optimized stimulus given an MNIST target image as the input.
Neural network. For the non-invertible neural network models, we use a feed-forward architecture consisting of three convolutional layers, each followed by batch normalization [26] and ELU activation [11]. The network is bookended by two linear layers at the input and output stages. Similar to the linear model, the feed-forward neural network is trained to map percepts to stimuli.
Invertible neural network with MMD. We use the open-source library FrEIA [1] for the implementation of the invertible network. Each INN layer is formed by a random permutation layer, Glow coupling block [1, 29], and an ActNorm layer [29]. The subfunction inside the coupling block contains two ReLU-activated linear layers with a hidden dimension of 512 and outputs both $s$ and $t$ simultaneously (Figure 4). A stack of 9 INN layers is empirically found to be sufficient and no additional performance gains are observed beyond that. The affine coupling layer [14, 29] is implemented using GLOWCouplingBlock in [1]. In contrast to AllInOneBlock in [1], GLOWCouplingBlock does not include ActNorm or invertible $1\times 1$ convolutions [29]. We also attempted to alter the dimensionality of the latent concatenations (Gaussian distribution in the forward path and zero paddings in the reverse path) either on the input or output side and applied the MMD loss in INNs, which did not result in performance improvement. During the training, we try to define either stimulus-to-percept or percept-to-stimulus as the forward process. We compare the MSE loss of both mappings and find that defining $h$ (Equation 1) as the percept-to-stimulus mapping makes its inverse function $h^{-1}$ easier to train, as shown in Figure 5, where the MSE losses are 0.2001 vs. 0.2756 at the step of 20000. Therefore, we report the results based on this setup in Table 2.
Conditional invertible neural network. The cINN architecture follows the design of the INN, with the additional conditioning to the subfunction at each layer [4, 2]. The percept or the MNIST target is concatenated as the conditioning during the training or the inference, respectively (Figure 3(b)). The conditional input $\mathbf{c}$, which is $\mathbf{y_{p}}$ during the training and $\tilde{\mathbf{x}}$ during the evaluation, is first transformed with a learnable linear layer to a size of $9\times 9$ before being passed to the subfunction. As introduced in Section 3.3, the cINN is trained in an unsupervised manner only with the NLL loss (Equation 7).
Table 2: Reconstruction quality evaluation from the predicted percepts to the original visual signals based on the optimized stimuli. Five approaches are compared with two resolutions (Res.) of the original MNIST input image of $9\times 9$ and $28\times 28$ pixels. Five evaluation metrics are reported, namely Mean Absolute Error (MAE), Mean Squared Error (MSE), Structural Similarity Index Measure (SSIM), Peak Signal-to-Noise Ratio (PSNR), and classification accuracy with a pre-trained classifier (ACC). The number of the parameters of each model is listed. The top results are marked in blue, while the second-best results are marked in red.
 Methods    Res. MAE $\downarrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ ACC $\uparrow$ #Param $\downarrow$
down 0.1682 0.0619 0.7248 30.4149 0.4272 1
linear 9 0.1657 0.0584 0.7246 30.6877 0.5608 6,561
NN $\times$ 0.1387 0.0436 0.7799 31.9620 0.5572 946,512
INN 9 0.1785 0.0599 0.7088 30.5656 0.4892 1,866,270
cINN 0.1373 0.0570 0.7651 30.8940 0.5480 1,936,710
down 0.3432 0.2040 0.3784 25.0973 0.2884 1
linear 28 0.3279 0.1999 0.4770 25.1453 0.5884 63,504
NN $\times$ 0.3340 0.1966 0.4901 25.2400 0.6192 4,270,900
INN 28 0.3231 0.1947 0.4937 25.2729 0.5764 18,063,390
cINN 0.3136 0.1845 0.4943 25.5161 0.5932 2,449,197
(a) Examples of random samples from $9\times 9$ stimuli to both $9\times 9$ and $28\times 28$ percepts using Axon Map Model [8] implemented in pulse2percept [7].
(b) Visualization of the reconstruction quality with five approaches. (I) denotes the original MNIST targets. (II) denotes the predicted percepts of $28\times 28$ pixels based on (III) optimized stimuli of $9\times 9$ electrodes.
Figure 7: Visualization during the training and the inference.
4.4 Likelihood and MSE
We illustrate whether the maximum likelihood training in the cINN approach leads to decrease the MSE loss, as it is reported that good likelihood does not always guarantee good samples in generative models [52]. The likelihood of the optimized stimulus depends on the likelihood of the latent $\mathbf{z}$ (Equation 3). Figure 6(a) shows the relationship between the log-likelihood of optimized stimulus (with $\mathbf{z}=\mathbf{0}$) and the MSE between the predicted percept and MNIST target (Figure 3(b)). For both random and MNIST samples, the correlation between the log-likelihood and the MSE can be observed. While setting $\mathbf{z}=\mathbf{0}$ should already give a good result, we further investigate if better $\mathbf{z}\neq\mathbf{0}$ can be found to yield higher likelihood and thus lower MSE. We select the worst 50 MNIST examples with the highest MSE and modify the latent $\mathbf{z}$ to maximize the total likelihood of optimized stimulus using gradient descent. Figure 6(b) shows that the MSE values are indeed decreased for most of the samples. Further closed-loop optimization w.r.t. $\mathbf{z}$ can be addressed in future work (e.g., [24, 37]).
4.5 Reconstruction Quality
The visualization of the reconstruction quality with a percept resolution of $28\times 28$ pixels are illustrated in Figure 7. We argue that negative stimulations (in red) can counteract the surrounding or elongated phosphenes, which is not applicable in downsampling. The predicted percepts (II) from the cINN show clearer turnings (e.g., the bottom left part in the image of the digit ’2’) and cleaner corners of the visual field. The advantages can be confirmed by quantitative assessments. Five evaluation metrics are reported in Table 2, namely Mean Absolute Error (MAE), Mean Squared Error (MSE), Structural Similarity Index Measure (SSIM) [58], Peak Signal-to-Noise Ratio (PSNR), and classification accuracy with a pre-trained classifier (ACC) [19, 62]. These metrics encompass pixel-level measurements (MAE, MSE), fidelity-oriented assessments (SSIM, PSNR), and semantic evaluations (ACC). The classifier is pre-trained on the MNIST dataset [32], achieving a validation accuracy of 0.9921 for downsampled $9\times 9$ pixel images and 0.9954 for the original $28\times 28$ pixel images. As shown in Table 2, cINNs demonstrate strong performance in both input resolutions, balancing pixel-level reconstruction with image-level semantic preservation.
4.6 Resolution Discrepancies
We experiment with two different visual signal resolutions ($9\times 9$ and $28\times 28$ pixels) using a fixed number of electrodes ($9\times 9$). Resolution discrepancies between the stimulus and the percept indicate that retinal prosthetic optimization is fundamentally a non-bijective inverse problem. With an increased percept resolution, cINN outperforms other approaches in various metrics. The number of the parameters of each model is listed in Table 2. The supervised models (linear model, feed-forward neural network and the MSE part in invertible neural networks) are susceptible to a higher resolution due to the fully-connected layers in the architectures, while this dense layer is only used to concatenate the target or percept to the subfunctions in the cINN approach. Consequently, the parameter count increases by only 1.26 times when using cINNs, compared to a 9.68-fold increase when using INNs with a corresponding 9.68-fold increase in the input size (Table 2). Thanks to their inherent invertibility and generativity, both INN-MMD and cINN show robust results when the input resolution is approximately 10 times larger than the electrode array of the retinal implant. Although deep neural networks have a large number of parameters, their implementation should not be problematic on modern portable hardware, potentially leading to substantial performance improvements of reconstruction quality.
5 Discussion
5.1 Limitations
Our work relies on a pre-defined Axon Map Model with fixed $\rho$ and $\lambda$ for simulation. However, these parameters are largely patient- and implant-dependent [9] and not facile to determine in real-world scenarios (cf. [16, 18]). Additionally, the pre-computed random samples, as illustrated in Figure 6(a), are not sufficiently general, making the study of more complex datasets a challenge. Finally, we did not compare our quantitative results with those of other papers in their original forms, as they are based on different assumptions and setups, such as the computational model used in retinal prosthetic simulation and the number of electrodes in the retinal implant. Nevertheless, we contend that our experiments include a diverse range of approaches, demonstrating the proof-of-concept for leveraging INNs in retinal prosthetic optimization.
5.2 Outlook
We anticipate our future studies to advance in three directions. First, supervised training of INN [3] can be further explored using different invertible architecture designs and training strategies. Second, simulation experiments can be conducted on more complex datasets utilizing other computational models to optimize stimuli. Third, eye movements of human beings, such as fixations, saccades, and micro-saccades, can be incorporated into the simulation and optimization processes.
6 Conclusion
We propose a stimulation optimization approach using conditional invertible neural networks for epiretinal implants and compare the reconstruction quality among downsampling, linear model, non-invertible and invertible neural networks. The cINN-based results outperform others w.r.t. various measures, especially with a higher resolution of visual inputs and percepts. Our findings highlight the potential application with this generative, invertible, and conditional neural network to enhance the quality of retinal prostheses.
Acknowledgements
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – grant 424556709/GRK2610.